# Collaborative Document Editing with Multiple Users and AI Agents

### v1

@@ 3.3.4. Task history in agent profiles
Whenever an agent executes a task, a task log entry is added to that agent’s profile (Figure 2F, G). This design supports transparency and scrutability in linking agents and tasks, and aligns with Park and Lee (2023), who found that co-writers value rationales for edits.
Figure 4. Example interaction flow when involving an AI agent via the comment UI: The user selects text and clicks the comment button in the floating toolbar (Figure 3D) to open a new comment (A). By typing “@ai”, the user brings up matching agent (or user) names (B); the user selects the default agent (“@aiauthor”) and enters a request (C). An indicator shows the agent is “typing” (D) while the LLM generates a response, which then appears in the thread (E). Clicking “Append” inserts the coloured text on the page (F), which the user edits before clicking “Approve” to finalize the change and close the comment. See Figure 5 for further examples.
[…]
@@ 3.4. Collaborative Comments for Users & Agents
Collaboration requires coordination and communication. As early as the 1990s, Sharples et al. (1993) highlighted communication as a core issue in collaborative writing software, and Rimmershaw (1992) noted the value of “meta-comments” as annotations tied to text, including author attribution. Our comment design builds on these insights and established conventions (e.g. Google Docs, Microsoft Word): Users select text and click the comment button in the floating toolbar (Figure 3D).
@@ 3.4.1. Agents respond via comments
Related work (Birnholtz and Ibara, 2012; Birnholtz et al., 2013) shows that edits carry social meaning and recommends presenting revisions as possibilities rather than definitive changes, especially when co-writers do not know each other well. This motivates our choice to have agents respond only through comments, never by editing text directly. Agents select text parts to comment on via a prompt-based procedure (Section 4). Our comment UI (Figure 4) supports this workflow with buttons for users to append a suggestion or replace selected text, plus a side-by-side view of original and suggested text (Figure 5A). Once an option is chosen, an “approve” button finalizes the change and closes the comment.
@@ 3.4.2. Referencing others and AI in comments
Users can reference co-writers in comments via “@” + user name, or agents via “@” + agent name. A generic default agent (“@aiAuthor”) is also available, allowing users to seek support without first creating a bespoke agent (Figure 4A-D). The referenced agent replies in the comment thread (Figure 4D, E), which co-writers can continue as usual (Figure 5B). Further agents can also be referenced, enabling them to “converse” with each other.
@@ 3.4.3. Ad-hoc tasks in comments
Referencing agents also enables writers to delegate something directly in a comment (Figure 4), bypassing task creation and the task list. This mirrors the common use of comments for delegating tasks to co-writers (Bødker et al., 2024), here extended to AI agents.
@@ 4. System Implementation
Our system consists of a backend and frontend application and connects to an LLM provider, as described in this section.
@@ 4.1. Frontend Application
Our frontend is based on React with React Bootstrap for styling and React Quill for the text editor. We developed a custom Quill plugin for creating annotations. Real-time collaboration and syncing is implemented with Yjs, a high-performance conflict-free replicated data type (CRDT) (Contributors, 2025). We used RTK Query for data fetching and local state management.
@@ 4.2. Backend Application
Our backend consists of two servers: one responsible for managing data, authentication, and authorization, and another dedicated to performing computationally intensive operations.
The main entry point is a Python server built with FastAPI, which launches Yjs and Socket.IO websockets to manage real-time data updates. The second server implements a Celery task queue to perform expensive background operations (i.e. agents’ document tasks). The servers connect to a MongoDB to persist and access data (e.g. documents, texts, comments, agent profiles).
@@ 4.3. AI Integration
Both FastAPI and Celery servers interact with an LLM through API calls. Our prototype in the study used OpenAI’s “gpt-4o-mini”. This is integrated into the backend via the openai and autogen libraries. The first is used for simple prompt-based queries, whereas the second handles the multi-turn conversations and agent-based interactions. The LLM is employed in the following cases:
@@ 4.3.1. Comment response generation
When users mention agents in a comment, an autogen group chat is created. The tagged agents are attached to the conversation and initialized with their profile information. Moreover, document text and goal, selected text, and previous conversation history for the selected text (if present) are added to the chat. The agent initialization prompt is shown in Appendix LABEL:lst:agent_prompt.
@@ 4.3.2. Agent CV suggestions and summary generation
For the suggestion feature on the CV page of the agent profile (Section 3.2.3), autogen is queried using a predefined prompt (Appendix LABEL:lst:agent_suggestions_prompt). Moreover, each time an agent is saved, its summary is updated via the autogen agent, using the prompt in Appendix LABEL:lst:agent_summary_prompt.
@@ 4.3.3. Document task title generation and automatic assignee selection
When a document task is saved, a title is generated based on its description using the openai API (prompt in Appendix LABEL:lst:task_title_prompt).
If the user does not specify an assignee, the system queries openai (prompt in Appendix LABEL:lst:task_assignee_prompt) to select the assignee based on the task and agent descriptions. The LLM returns a confidence score; if the confidence is 85% or higher, the recommended agent is assigned. Otherwise, the system assigns the default “AI Author” agent.
@@ 4.3.4. Task execution across the document
When a task is executed on the entire document, the Yjs websocket server sends the task to the Celery server. Its response is then applied to the Yjs document and stored in the database. These are the steps in detail:
@@
    Generating selection proposals (Celery): The first agent assignee selects relevant text segments. Based on the assignee and task description, the LLM is queried to return the text segments, their length, reasoning, and confidence score (prompt in Appendix LABEL:lst:task_text_selection_prompt).
@@
    Generating responses (Celery): For each valid segment, a conversation with the task assignees is initiated using the same process as described in Section 4.3.1. The results are returned to the Yjs websocket server.
@@
    Integrating responses (Yjs websocket): New annotations and comments are integrated into the Yjs document, while existing annotations are updated. The document is updated in the database.
@@ 4.4. Autonomous tasks
To detect autonomous triggers and execute tasks, the Yjs websocket server continuously monitors document activity and evaluates the trigger conditions. Once a condition is met, it requests task execution from the Celery server. After receiving the result, it applies the response to the document. Table 1 gives an overview of the triggers with implementation notes.
@@
[…]
A two-column table listing triggers for autonomous tasks and the corresponding conditions that cause each trigger to fire. Trigger condition Implementation In short intervals As soon as someone enters the document socket room, schedule a job to run every 5 minutes. Skip if a job is already scheduled. After some time of inactivity of all users Debounce 2 minutes after the last detected user activity (text or comment updates). When all users are offline When the last online user disconnects from the document socket room. Once document is saved When a user clicks the Save button. After several collaborative edits The Yjs document includes a shared array that tracks the users who typed in the document. If the list has more than two users, it indicates a collaborative edit. The array is cleared after the trigger is executed to detect future edits.
@@ 5. User Study
We conducted an exploratory user study with our prototype as a design probe to investigate how co-writers create AI agents and interact with them through our design features. The study consisted of a collaborative writing task using the prototype, questionnaires, and a semi-structured group interview at the end. Each group had one week to work on the task.
@@ 5.1. Participants
We recruited 30 participants (14 female and 16 male) in 14 groups by sharing a study description and registration form in our network across a few universities, where we reached mostly faculty and students. Ten groups signed up as a group directly, three groups signed up individually but knew each other, and one group (7) was formed by us from two individual study registrations.
Participants’ age ranged from 22 to 56 years (median 27), and they all had sufficient knowledge of English. Their median frequency of writing documents was daily and the frequency of collaborative writing was weekly. The median use of LLMs was also daily. Their writing domains included academic study, professional work, research, and journalism. Participants were compensated with $25\text{\,}\text{\texteuro}$.
See Table 3 in the Appendix for further details.
@@ 5.2. Study Procedure
The procedure was coordinated via the StudyAlign web application (Lehmann and Buschek, 2025), which integrated the questionnaires, writing task and application description page, and access to the prototype. Once participants received the study link and provided informed consent, they were free to work with the prototype for a week. We recorded interactions with the prototype. The semi-structured group interviews were conducted using Microsoft Teams and were audio-recorded and transcribed. The study followed our institute’s regulations (e.g. information on goals, process, data protection, opt-out and consent form). We next describe each step in more detail.
@@ 5.2.1. Introductory meeting (15 minutes)
Based on feedback from the first three groups, we concluded that a quick tour of the design features would help the participants get started with the prototype. Thus, we conducted introductory online meetings for the remaining groups, where one researcher demonstrated the features via screen sharing.
@@ 5.2.2. Writing with the prototype (7 days)
For this main part of the study, our web app guided participants through the following steps:
@@
    Task briefing (20 minutes): Participants could choose between two collaborative writing tasks: either working on a real task of their own, or, if none was available, writing an essay about the positive and negative aspects of AI in daily life. The briefing also included a page explaining the prototype’s features through text and videos.
@@
    Writing task (7 days): Participants used a join code to access their shared document. They could then collaborate on the task. Each group was free to decide how to approach the task and collaborate on the document.
@@
    Final questionnaire (25 minutes): Participants rated their experience with the prototype and answered SUS (Aaron Bangor and Miller, 2008) and CSI (Carroll et al., 2009) questionnaires. They also evaluated the individual features, and provided open comments on their negative and positive aspects.
@@ 5.2.3. Semi-structured interview (30 minutes)
The last step with each group was a video call where we interviewed the group, including questions on likes/dislikes, collaboration, challenges, agents, tasks, and triggers. We chose this group format so that team members could reflect on and discuss their collaborative experience together.
@@ 6. Results
6.1. Collaboration and Tool Use
In our initial questionnaire, we asked participants to rate how often they work on texts and how they collaborate (see Table 3 in the Appendix). Further tools that participants use when collaborating on text documents included chat tools like Slack, Mattermost, and Teams (see Figure 10 in the Appendix). Sixteen participants use the chat tools daily, six weekly, four monthly, and four less than monthly.
6.2. Usability Ratings
In the final questionnaire, we assessed the system usability scale (SUS) and the creativity support Index (CSI). Ratings on the SUS score had a mean score of 66.92 (SD=12.96, min=45, max=90), see Figure 7 in the Appendix. This score is slightly below the average benchmark of 68. The CSI score yielded a mean of 67.21 (SD=15.60, min=31, max=91). As shown in Table 4 in the Appendix, the most valued factors of the CSI were “Enjoyment” and “Collaboration”.
In addition to these standardised questionnaires, we asked own Likert items on general feedback (Figure 8) and the AI agents (Figure 9). Most users found the interaction with the system natural and felt it helped them complete tasks more efficiently. Around a quarter of participants (23.33%) did not feel in control of the text. The agent features were rated positively overall but participants were divided about if one agent is enough or not.
6.3. Interaction Behaviour
We logged how users interacted with the agent, task, and comment features. One group accidentally accessed the prototype outside of our study setup, such that logs were not recorded. We reconstructed their interactions from database data where possible and note results where this group (G4) is excluded from the analysis.
@@ 6.3.1. Metrics on Agent Creation
In total, 39 custom agents were created with an average of 2.79 per group (SD=1.25, min=0, max=4). One group did not create a custom agent. 38.46% of agents were created in the first quartile of each group’s normalised editing time, 28.2% were created in the second and third quartiles, and the remaining 33.33% were created in the last quartile. This behaviour reflects initial exploration as well as a demand for agents for drafting and finalising text.
The following results on agent usage exclude G4. Agents could be defined through the “CV” and notes features (Figure 2). A total of 54 CV sections were created (28 skill sections and 26 expertise sections). Participants manually added 71 values (33 skills, 38 expertises). Responding to participants’ requests for suggestions, the system suggested 174 values, and participants added 45.4% (79) of them (36 skills, 43 expertises). Notes were added to custom agents in four cases only. Another three notes were added to agents extended from presets. Thus, people mainly used the structured UI (“CV”) for defining agents.
Once created, agents were updated in eleven cases; ten such edits happened by the agent’s respective creator, and only one edit was by another team member. This indicates that agents were seen as belonging to their creator. The qualitative results shed more light on this (Section 6.4.1).
@@ 6.3.2. Metrics on Tasks from the Task List
Participants created a total of 67 tasks via the task list feature (Figure 3), with an average of 5.15 per group (SD=2.19, min=1, max=9). The following results exclude G4. Most such tasks (88.06%) were created by the author of the agent to which the task was assigned. The tasks from the task list were triggered manually 113 times; 84.7% by the task’s creator. Task editing occurred 10 times. We also analysed task creation across the normalised editing time of each group. Most tasks were created in the last quartile (41.67%) and in the third quartile (23.33%). The remaining task creations occurred in the first (21.66%) and second (13.33%) quartile.
@@ 6.3.3. Metrics on Comments
We counted a total of 468 comments; 376 were added by agents in response to running tasks, while participants added 92 comments. G2, G5 and G12 never used comments for ad-hoc task delegation to AI. The other eleven groups created a mean of 8.36 comments (SD=7.19, min=1, max=22). We analysed specific discussion patterns considering comments and their replies, as shown in Table 2. These reveal that only short discussions happened in comments, for example single turn human-AI discussions. Note that most groups reported to use further communication tools, explaining the small number of human-human discussions in comments in our study.
@@
A three-column table listing the five discussion patterns that appeared in the log data. Each pattern is described and has a count. Discussion pattern Description Count human-human comment by user, reply by team members 4 human-AI comment by user, AI replies; single-turn 57 human-AI negotiation comment by user, AI replies, human replies; multi-turn 17 human-team-AI comment by user, AI replies, team members reply; multi-turn 6 AI-human comment by AI, at least one user replies 7
@@
To accept agents’ suggestions in comments, we offered three methods: Appending text at the end of the selected text in the document, replacing the selected text, and copying the suggestion to the clipboard. Excluding G4 again, suggestions were appended 76 times; 50 times by the comment’s creator and 26 times by others (19 times when the comment was created by a user, 7 times when it was created by AI). The replace method was used 60 times; 29 times by the comment’s creator and 31 times by others (23 times for user-created comments, 8 times for AI-created ones). The copy to clipboard method was used less frequently (16 times): 9 times by the comment’s creator, and 7 times by others (6 times for user-created comments, once for an AI-created comment). This use of the three accept methods shows that agent-initiated suggestions were less often accepted than user-initiated suggestions.
@@
After normalising the editing time of each group, we analysed when participants posted comments. They posted 4.83% of comments in the first quartile, 21.38% in the second, 11.03% in the third, and 62.76% in the last quartile. This behaviour indicates that participants sought support from agents after writing a first draft, particularly in order to complete it.

### v2

@@ 3.3.4. Task history in agent profiles
Whenever an agent executes a task, a log entry is added to that agent’s profile (Fig. 2F, G). We added this for transparency (i.e. all users can see all agents and profiles) and scrutability in linking agents and tasks (i.e. all users can inspect all task runs by all agents), in line with Park and Lee (2023), who found that co-writers value rationales for edits.
Figure 4. Example interaction flow when involving an AI agent via the comment UI: The user selects text and clicks the comment button in the floating toolbar (Fig. 3D) to open a new comment (A). By typing “@ai”, the user brings up matching agent (or user) names (B); the user selects the default agent (“@aiauthor”) and enters a request (C). An indicator shows the agent is “typing” (D) while the LLM generates a response, which then appears in the thread (E). Clicking “Append” inserts the coloured text on the page (F), which the user edits before clicking “Approve” to finalize the change and close the comment. See Fig. 5 for further examples.
@@ 3.4. Collaborative Comments for Users & Agents
Already in the 1990s, Sharples et al. (1993) highlighted communication as a core issue in collaborative writing software, and Rimmershaw (1992) noted the value of “meta-comments” as annotations tied to text. Our comment design builds on these insights and established conventions (e.g. Google Docs, Microsoft Word): Users select text and click the comment button in the floating toolbar (Fig. 3D).
@@ 3.4.1. Agents respond via comments
Related work (Birnholtz and Ibara, 2012; Birnholtz et al., 2013) shows that edits carry social meaning and recommends presenting revisions as possibilities rather than definitive changes, especially when co-writers do not know each other well. This motivates our choice to have agents respond through comments, not by editing text directly. Agents select text parts to comment on via a prompt-based procedure (Section 4). Our comment UI (Fig. 4) provides buttons for users to append a suggestion or replace selected text, plus a side-by-side view of the original and the suggestion (Fig. 5A). Once an option is chosen, an “approve” button finalizes the change and closes the comment.
@@ 3.4.2. Referencing others and AI in comments
Users can reference co-writers in comments via “@” + user name, or agents via “@” + agent name. A generic default agent (“@aiAuthor”) provides support without first creating a bespoke one (Fig. 4A-D). The referenced agent replies in the comment thread (Fig. 4D, E), which co-writers can continue as usual (Fig. 5B). Further agents can also be referenced, to “converse” with each other.
@@ 3.4.3. Ad-hoc tasks in comments
Referencing agents also enables writers to delegate something directly in a comment (Fig. 4), bypassing task creation and task list. This mirrors using comments for delegating tasks to co-writers (Bødker et al., 2024).
@@ 4. System Implementation
Our system consists of a backend and frontend and connects to an LLM provider.
@@ 4.1. Frontend Application
Our frontend uses React with React Bootstrap and React Quill for the editor. We developed a custom Quill plugin for creating annotations. Real-time collaboration and syncing is implemented with Yjs, a conflict-free replicated data type (CRDT) (Contributors, 2025). We used RTK Query for data fetching and local state management.
@@ 4.2. Backend Application
Our Python server is built with FastAPI, launching Yjs and Socket.IO websockets to manage real-time data updates. A second server implements a Celery task queue for expensive background operations (i.e. agents’ document tasks). The servers connect to a MongoDB (e.g. persisting documents, texts, comments, agent profiles).
@@ 4.3. AI Integration
Both servers call an LLM API. Our prototype used OpenAI’s “gpt-4o-mini”. We chose this model to represent LLM capabilities in our study in line with a currently widely used product (ChatGPT), with fast response times. This is integrated into the backend via the openai and autogen libraries. The first is used for simple prompt queries, the second handles multi-turn conversations and agent-based interactions. The LLM is employed in the following cases:
@@ 4.3.1. Comment response generation
When users mention agents in a comment, an autogen group chat is created. The tagged agents are attached to the conversation and initialized with their profile (prompt in LABEL:lst:agent_prompt). Document text and goal, selected text, and conversation history for the selected text (if present) are added.
@@ 4.3.2. Agent CV suggestions and summary generation
For CV suggestions (Section 3.2.3), autogen is queried using the prompt in LABEL:lst:agent_suggestions_prompt. Each time an agent is saved, its summary is updated via the autogen agent, using the prompt in LABEL:lst:agent_summary_prompt.
@@ 4.3.3. Document task title generation and automatic assignee selection
When a document task is saved, a title is generated based on its description using the openai API (prompt in LABEL:lst:task_title_prompt). If the user does not specify an assignee, the system queries openai (prompt in LABEL:lst:task_assignee_prompt) to select one based on task and agent descriptions. The LLM returns a confidence score; if it is 85% or higher, the recommended agent is assigned. Otherwise, the system assigns the default agent.
@@ 4.3.4. Task execution across the document
When a task is executed on the entire document, the Yjs websocket server sends the task to the Celery server. Its response is then applied to the Yjs document and stored in the database. In detail:
@@
    Generating selection proposals (Celery): The first agent assignee selects relevant text segments. Based on the assignee and task description, the LLM is queried to return the segments, their length, reasoning, and confidence score (prompt in LABEL:lst:task_text_selection_prompt).
@@
    Generating responses (Celery): For each valid segment, a conversation with the task assignees is initiated using the process as in Section 4.3.1. Results are returned to the Yjs websocket server.
@@
    Integrating responses (Yjs websocket): New annotations and comments are integrated into the Yjs document, while existing annotations are updated.
@@ 4.4. Autonomous Tasks
The Yjs websocket server continuously monitors document activity and evaluates the trigger conditions (Table 2), to request task execution from the Celery server.
@@
A two-column table listing triggers for autonomous tasks and the corresponding conditions that cause each trigger to fire.
Trigger condition Implementation
In short intervals As soon as someone enters the document socket room, schedule a job to run every 5 minutes. Skip if a job is already scheduled.
After some time of inactivity of all users Debounce 2 minutes after the last detected user activity (text or comment updates).
When all users are offline When the last online user disconnects from the document socket room.
Once document is saved When a user clicks the Save button.
After several collaborative edits The Yjs document includes a shared array that tracks the users who typed in the document. If the list has more than two users, it indicates a collaborative edit. The array is cleared after the trigger is executed to detect future edits.
@@ 5. User Study
We conducted an exploratory user study with our prototype as a technology probe to investigate how co-writers collaboratively create and interact with AI agents. We motivate our approach by following the mapping of the three goals of technology probes (Hutchinson et al., 2003) to LLM-involved design research by Feng et al. (2025). A main interest with our study maps to the social science goal: understanding how teams experience the integration of (customisable) AI agents into their shared documents. Another interest maps to the design goal: “inspiring users and researchers to think about new technologies” (Hutchinson et al., 2003), in our case collaborative integration of agents in shared documents. In addition, our probe serves the engineering goal of field-testing, for a prototype release to facilitate future studies.
@@ 5.1. Participants
We recruited 30 participants (14 women, 16 men) in 14 groups by sharing a study call in our network across a few universities, where we reached mostly faculty and students (Table 4 in Appendix). Ten groups signed up directly, three signed up individually but knew each other, and one group (G7) was formed by us from two individual registrations. Participants’ age ranged from 22 to 56 years (median 27), and they all had sufficient knowledge of English. Their median frequencies of using LLMs and writing documents was daily, and collaborative writing was weekly. Their writing domains included academic study, professional work, research, and journalism. They were compensated with $25\text{\,}\text{\texteuro}$.
@@ 5.2. Study Procedure
Once participants received the study link and provided informed consent, they were free to work with the prototype for a week. We recorded interactions and conducted semi-structured group interviews using Microsoft Teams, which were audio-recorded and transcribed. The study followed our institute’s regulations (e.g. information on goals, process, data protection, opt-out and consent form).
@@ 5.2.1. Introductory meeting (15 minutes)
Based on feedback from the first three groups, we concluded that a quick tour of the features would be helpful. Thus, we conducted online introductions for the remaining groups, where one researcher demonstrated the features via screen sharing.
@@ 5.2.2. Writing with the prototype (7 days)
The procedure was coordinated via the StudyAlign web application (Lehmann and Buschek, 2025), which integrated the questionnaires, writing task and description page, and access to the prototype.
@@
    Task briefing (20 minutes): Participants chose between two tasks – either working on a real task of their own, or, if none was available, writing an essay on positive and negative aspects of AI in daily life. The briefing explained the prototype’s features through text and videos.
@@
    Writing task (7 days): Participants used a join code to access their shared document. Groups were free to decide how to approach the task and collaboration.
@@
    Final questionnaire (25 minutes): Participants answered SUS (Aaron Bangor and Miller, 2008) and CSI (Carroll et al., 2009) questionnaires, evaluated individual features, and provided comments on pros and cons.
@@ 5.2.3. Semi-structured interview (30 minutes)
Group interviews via video calls included questions on likes/dislikes, collaboration, challenges, agents, tasks, and triggers. This group format allowed teams to reflect on their experience together.
@@ 6. Results
We first report results from logging data and questionnaires (Section 6.1) as a summary of interaction in the study, before turning to our main qualitative analysis (Section 6.2).
6.1. Usability and Interaction Behaviour
While our focus is on qualitative exploration, we also logged interactions to support and contextualise the qualitative data. One group (G4) accidentally accessed the prototype outside of our study setup, such that logs were not recorded. We reconstructed their interactions from our database where possible and note where they are excluded.
6.1.1. Usability ratings
To provide basic background information, participants initially rated how often they work on texts and how they collaborate (Table 4 in the Appendix). Further tools used when collaborating on documents included Slack, Mattermost, and Teams (Fig. 10 in the Appendix). Sixteen use such chat tools daily, six weekly, four monthly, and four less than monthly.
We also assessed usability measures as an overview of participants’ perception of the design features, which we enrich per feature in our qualitative analysis in Section 6.2.5. Ratings on the system usability scale (SUS) had a mean score of 66.92 (SD=12.96, min=45, max=90), see Fig. 7 in the Appendix. This score is slightly below the average benchmark of 68. The CSI yielded a mean of 67.21 (SD=15.60, min=31, max=91). As shown in Table 5 in the Appendix, the most valued factors of the CSI were “Enjoyment” and “Collaboration”. We further asked Likert items on general feedback (Fig. 8) and agents (Fig. 9). Most found the interaction natural and felt it helped them reach their goals faster. Around a quarter (23.33%) did not feel in control of the text. The agent features were rated positively overall but participants were divided about whether one agent is enough or not.
@@ 6.1.2. Metrics on agent creation
Participants created 39 agents, 2.79 per group (SD=1.25, min=0, max=4). 38.46% were created in the first quartile of each group’s normalised editing time, 28.2% in the second and third, and 33.33% in the last. This reveals: (1) initial exploration, typical for experiencing new features – yet not exclusively, suggesting (2) emerging needs or preferences addressed with agents created later. A variety of needs is also reflected in the text analyses of profiles and tasks in Section 6.3.
The following results exclude G4. Agents could be defined through the “CV” and notes features (Fig. 2): 54 CV sections were created (28 skill sections and 26 expertise sections) with 71 values added manually (33 skills, 38 expertises). Responding to participants’ requests for suggestions, the system suggested 174 values, and participants added 45.4% (79) of those (36 skills, 43 expertises). Notes were added to custom agents in four cases only. Another three notes were added to agents extended from presets. These results show that people mainly used the structured UI (“CV”) for defining agents. In contrast, the free-form notes were not seen as adding further value.
Once created, agents were updated in eleven cases; ten times by the agent’s respective creator, and only once by another team member. Together with the qualitative results (Section 6.2.1), this shows that agents were seen as belonging to their creator.
@@ 6.1.3. Metrics on tasks from the task list
Participants created 67 tasks via the task list feature (Fig. 3), 5.15 per group (SD=2.19, min=1, max=9), excluding G4. Most (88.06%) were created by the author of the agent to which the task was assigned. The tasks from the task list were triggered manually 113 times; 84.7% by the task’s creator. Task editing occurred 10 times. For tasks with autonomous triggers (7 tasks), the most common trigger was “After some time of inactivity of all users” (3 tasks), followed by “In short intervals” (2), and “Once document is saved” and “When all users are offline” (1 each).
We also analysed task creation across the normalised editing time per group. Most tasks were created in the last quartile (41.67%) and in the third (23.33%), compared to the first (21.66%) and second (13.33%). Combined with the agent results, this indicates that custom agents are created before custom tasks. Participants’ rationales for valuing custom agents explain this: Some saw agent profiles as a way to “preload” context for later task prompts (Section 6.2.3).
@@ 6.1.4. Metrics on comments
We counted 468 comments; 376 were added by agents in response to running tasks, while participants added 92 comments. G2, G5 and G12 never used comments for ad-hoc task delegation to AI. The other groups created a mean of 8.36 comments (SD=7.19, min=1, max=22). Table 3 shows specific discussion patterns, revealing that only short discussions happened in comments. For contextualisation, note that most groups reported to use further communication tools, explaining the small number of human-human discussions via comments here.
@@
A four-column table listing the five discussion patterns that appeared in the log data. Each pattern is described and has a count.
Discussion pattern Description How this was used Count
human-human comment by user, reply by team members coordinating tasks 4
human-AI comment by user, AI replies; single-turn delegate task; directly accept/reject AI suggestion 57
human-AI negotiation comment by user, AI replies, human replies; multi-turn refinement of AI suggestions 17
human-team-AI comment by user, AI replies, team members reply; multi-turn work together on goals (e.g. brainstorming); involve different agents 6
AI-human comment by AI, at least one user replies refinement of AI initiated suggestions 6
@@
To accept agents’ suggestions in comments, we offered three methods: Appending text at the end of the selected text, replacing the selected text, and copying the suggestion to the clipboard. Excluding G4 again, suggestions were appended 76 times; 50 times by the comment’s creator, 26 times by others (19 times when the comment was created by a user, 7 times when it was created by AI). Replace was used 60 times; 29 times by the comment’s creator, 31 times by others (23 times for user-created comments, 8 times for AI). Copy to clipboard was only used 16 times: 9 times by the comment’s creator, 7 times by others (6 times for user-created comments, once for AI). Thus, agent-initiated suggestions were less often accepted than user-initiated ones. This had two reasons: (1) User-initiated AI suggestions reply to a user’s comment on selected text and thus respond at the user’s current focus, making acceptance more likely (“as is” or for further editing them). (2) In contrast, agent-initiated comments from running a document task can result in a high number of comments (anywhere); when feeling overwhelmed, users ignored some of them. See Section 6.2.5 for details on this from the qualitative insights, and Section 6.1.5 for further related logging analyses.
@@
We analysed when participants posted comments: 4.83% of comments were posted in the first quartile, 21.38% in the second, 11.03% in the third, and 62.76% in the last. This indicates that participants typically sought support from agents after writing a first draft.
6.1.5. Metrics on collaboration
Here we provide two metrics on collaboration as contextualisations for the qualitative results on how teams collaborated in our study. These analyses exclude G4.
Interaction timing and engagement with (each others’) AI comments
We computed the times from AI adding a comment until a user accepting it. The grand mean was 19.98 minutes (SD=121.41, min=0.01, max=1363.53). Users accepted AI suggestions responding to human comments (136) after 22.26 minutes (SD=128.21, min=0.07, max=1363.53). They accepted AI suggestions from AI-initiated comments less often (16) but faster, after 0.67 minutes (SD=0.70, min=0.01, max=2.07). The difference is partly due to acceptances by group members with a long time gap, such as when working asynchronously. Concretely, users accepted 48 suggestions made by AI in response to another user’s comment, with a mean of 59.56 minutes (SD=212.06, min=0.14, max=1363.53). They accepted (88) AI suggestions in response to their own comments faster, with a mean of only 1.92 minutes (SD=5.08, min=0.07, max=34.61). Thus, AI comments became more temporary suggestion displays for a single user when integrated into the requesting user’s own workflow, whereas they remained (longer lasting) action items when left for others (cf. Section 6.2.4).
Synchronicity with each other and AI
We measured interaction switches to analyse how interleaved each group collaborated. We ordered input events chronologically and counted switches of author IDs per group. We also included agents via AI comment events and acceptance events. The grand mean was 219.07 interaction switches (SD=285.28, min=10, max=990). See Table 6 in the appendix for details – the numbers, both absolute and relative (switches to human vs AI), help further characterise our sample: We observed a range of collaboration styles, with some groups collaborating with highly intertwined edits (e.g. G1, G14), and others more asynchronously or in sequence (e.g. G8, G11). The same holds for working with the agents, with which some groups had more back-and-forth “handovers” than others. We see these results as another way of characterising the diversity of workflows in our study and focus on the qualitative data for further insights into these collaboration patterns (Section 6.2.4).